# QuantFactor REINFORCE: Mining Steady Formulaic Alpha Factors with Variance-bounded REINFORCE

### v1

@@ V Numerical Results
In this section, we compare the efficiency of the QFR algorithm with other typical state-of-the-art reinforcement learning algorithms in solving MDPs for mining formulaic alpha factors. Furthermore, we benchmark the proposed QFR against other existing factor mining algorithms in terms of the factors they generate and their performance in more realistic trading scenarios.
@@ V-A Environment Settings
The raw data sourced from the Chinese A-shares market, as well as the US stock market, specifically focusing on the constituent stocks of the China Securities Index 300 (CSI300), the China Securities Index 500 (CSI500), the China Securities Index 1000 (CSI1000), the S&P 500 Index (SPX), the Dow Jones Industrial Average (DJI), and the NASDAQ 100 Index (NDX) are utilized to model the MDPs for mining formulaic alphas in our experiment. To ensure reproducibility, we have only identified six primary features to generate the formulaic alphas, which include opening price (open), closing price (close), highest price (high), lowest price (low), trading volume (volume), and volume-weighted average price (vwap). Our objective is to generate formulaic alphas that exhibit a high IC with respect to the 5-day asset returns. The dataset is split into three subsets: a training set spanning from 2016/01/01 to 2020/01/01, a validation set spanning from 2020/01/01 to 2021/01/01, and a test set spanning from 2021/01/01 to 2024/01/01. All price and volume data have been forward-dividend-adjusted respected to the adjustment factors on 2023/01/15.
@@
In addition, both the QFR and AlphaGen are used for solving the MDPs defined in Section III-B. In order to demonstrate the state-of-the-art performance of QFR in this reinforcement learning task, some respected reinforcement learning baseline algorithms, including TRPO[54], PPO[44] and A3C[55] are also utilized in the experiment.
Figure 5: Performance of our method and state-of-the-art reinforcement learning methods for mining formulaic alpha factors in the constituent stocks of the China Securities Index 300 (CSI300), the China Securities Index 500 (CSI500), the China Securities Index 1000 (CSI1000), the S&P 500 Index (SPX), the Dow Jones Industrial Average (DJI), and the NASDAQ 100 Index (NDX). All the curves are averaged over 5 different random seeds, and half of the standard deviation is shown as a shaded region.
To demonstrate the effect caused by stochasticity in the training process, each experimental combination with an indeterministic training process is evaluated with 5 different random seeds. The hyperparameters of MLP, XGBoost and LightGBM are set according to the benchmarks given by Qlib. The hyperparameters of GP are set according to the gplearn framework. The actor network and critical network of PPO, A3C and TRPO share a base LSTM feature extractor, which has a 2-layer structure with a hidden layer dimension of 128. The drop out rate is set to 0.1. The separate value and policy heads are MLPs with two hidden layers of 64 dimensions. PPO clipping range $\epsilon$ is set to 0.2. TRPO trust region constraint upper bounds $\delta$ is set to $1.1\cdot 10^{-5}$. All the operator tokens used in our experiment are shown in Table III. The simulation is performed by a single machine with an Intel Core i9-13900KCPU and two NVIDIA GeForce RTX 4090 GPUs.
TABLE III: All the Operator Tokens Used in the Experiment
@@
Ref($x,t$) Time-Series
Mean($x,t$), Medium($x,t$) Sum($x,t$) Time-Series
Std($x,t$), Var($x,t$) Time-Series
Max($x,t$), Min($x,t$) Time-Series
Mad($x,t$) Time-Series
Delta($x,t$) Time-Series
WMA($x,t$), EMA($x,t$) Time-Series
Cov($x,y,t$), Corr(x,y,t) Time-Series
@@ V-B Comparisons with Other RLs
When solving the MDPs defined in Section III-B, QFR algorithm demonstrates superior performance compared to PPO that is adopted in AlphaGen[14]. We present experimental results on the constituent stocks of six indices in China and the United States using typical state-of-the-art RL algorithms, including TRPO, PPO, and A3C, as shown in Fig 5. Due to the fact that QFR utilizes reward shaping, while the baseline algorithm does not, reward is not used as a metric when comparing the learning curves of various algorithms. Instead, Rank Information Coefficient (Rank IC) is used to substitute for reward in evaluating the performance of the policy network during training. The Rank IC assesses the association between the ordinal positions of alpha values and the sequences of stock return rankings. Rank IC is just the IC of ranked data, defined as $RankIC\left(\mathbf{z}^{\prime}_{t},\mathbf{y}_{t}\right)=IC\left(r(\mathbf{z}^{\prime}_{t}),r(\mathbf{y}_{t}\right))$, where $r(\cdot)$ is the ranking operator. This metric is the higher the better. It can be observed that our method can outperform the other RL algorithms in all six indices, especially at the end of the training process. Compared to the best-performing PPO algorithm adopted by AlphaGen, our algorithm improves by 3.83% on the metric, which further validates the effectiveness of discarding elements related to the value network, introducing a subtractive baseline value, as well as the reasonable reward shaping.
V-C Factors Evaluation
@@
We applied the baseline algorithms and our QFR algorithm to the constituent stocks of the CSI300 and CSI500 indices, and evaluated their performance based on two metrics: IC and Rank IC. The IC is delineated in (1). Both of the metrics are the higher the better. Note that all the algorithm is only optimized against the IC metric. The results are shown in Table IV. The performance of MLP, XGBoost, and LightGBM is inferior due to the usage of the open-source set of alphas. On the other hand, AlphaGen tends to converge to local optima, leading to overfitting on the training data. Although GP mitigates this issue by preserving a diverse population, it still struggles to generate alphas that exhibit synergy in combination. The QFR algorithm excels in solving MDPs for mining formulaic alpha factors compared to AlphaGen. We attribute the superior performance of QFR to the acceleration of the convergence process by discarding elements related to the value network, as well as the reasonable reward shaping.
@@ V-D Investment Simulation
TABLE IV: Performance of Mined Factors on CSI300 and CSI500
@@ VI Conclusion
In this paper, we have proposed a novel reinforcement learning algorithm, QuantFactor REINFORCE (QFR), for mining formulaic alpha factors. QFR leverages the advantages of discarding the value network while addressing its limitations of high variance by introducing a greedy baseline. Additionally, the incorporation of IR as a reward shaping mechanism encourages the generation of stable alpha factors that can better adapt to changing market conditions.

### v2

@@ V Numerical Results
In this section, we numerically evaluate QFR, comparing it with both state-of-the-art RL algorithms and other factor mining methods. Our experimental study is comprised of six stock datasets (detailed in Section V-A), and evaluates the efficiency of various RL algorithms in solving MDPs for mining formulaic alpha factors in Section V-B, after which we study the hyper-parameters in reward shaping in Section V-C, and considers five factor mining methods (discussed in Section V-D and V-E). Finally, the ablation study confirms the importance of the two improvements in Section V-F.
@@ V-A Environment Settings
The raw data sourced from the Chinese A-shares market, as well as the US stock market, specifically focusing on the constituent stocks of the China Securities Index 300 (CSI300), the China Securities Index 500 (CSI500), the China Securities Index 1000 (CSI1000), the S&P 500 Index (SPX), the Dow Jones Industrial Average (DJI), and the NASDAQ 100 Index (NDX) are utilized to model the MDPs for mining formulaic alphas in our experiment. To ensure reproducibility, we have only identified six primary features to generate the formulaic alphas, which include opening price (open), closing price (close), highest price (high), lowest price (low), trading volume (volume), and volume-weighted average price (vwap). Our objective is to generate formulaic alphas that exhibit a high IC with respect to the ground-truth 5-day asset returns. The dataset is split into three subsets: a training set spanning from 2016/01/01 to 2020/01/01, a validation set spanning from 2020/01/01 to 2021/01/01, and a test set spanning from 2021/01/01 to 2024/01/01. All price and volume data have been forward-dividend-adjusted respected to the adjustment factors on 2023/01/15.
@@
In addition, both the QFR and AlphaGen are used for solving the MDPs defined in Section III-C. In order to demonstrate the state-of-the-art performance of QFR in this reinforcement learning task, some respected reinforcement learning baseline algorithms, including TRPO[60], PPO[21] and A3C[61] are also utilized in the experiment.
Figure 5: Performance of our method and state-of-the-art reinforcement learning methods for mining formulaic alpha factors in the constituent stocks of CSI300, CSI500, CSI1000, SPX, DJI, and NDX. All the curves are averaged over 5 different random seeds, and half of the standard deviation is shown as a shaded region.
To demonstrate the effect caused by stochasticity in the training process, each experimental combination with an indeterministic training process is evaluated with 5 different random seeds. The hyperparameters of MLP, XGBoost and LightGBM are set according to the benchmarks given by Qlib. The hyperparameters of GP are set according to the gplearn framework. The actor network and critical network of PPO, A3C and TRPO share a base LSTM feature extractor, which has a 2-layer structure with a hidden layer dimension of 128. The drop out rate is set to 0.1. The separate value and policy heads are MLPs with two hidden layers of 64 dimensions. PPO clipping range $\epsilon$ is set to 0.2. TRPO trust region constraint upper bounds $\delta$ is set to $1.1\cdot 10^{-5}$. All the operator tokens used in our experiment are shown in Table IV. The simulation is performed by a single machine with an Intel Core i9-13900KCPU and two NVIDIA GeForce RTX 4090 GPUs.
TABLE IV: All the Operator Tokens Used in the Experiment
@@
Ref($x,l$) Time-Series
Mean($x,l$), Medium($x,l$) Sum($x,l$) Time-Series
Std($x,l$), Var($x,l$) Time-Series
Max($x,l$), Min($x,l$) Time-Series
Mad($x,l$) Time-Series
Delta($x,l$) Time-Series
WMA($x,l$), EMA($x,l$) Time-Series
Cov($x,y,l$), Corr($x,y,l$) Time-Series
@@ V-B Comparisons with Other RLs
When solving the MDPs defined in Section III-C, QFR demonstrates superior performance compared to PPO that is adopted in AlphaGen[19]. We present experimental results on the constituent stocks of six indices in China and the United States using typical state-of-the-art RL algorithms, including TRPO, PPO, and A3C, as shown in Fig 5. Due to the fact that QFR utilizes reward shaping, while the baseline algorithm does not, reward is not used as a metric when comparing the learning curves of various algorithms. Instead, Rank Information Coefficient (Rank IC) is used to substitute for reward in evaluating the performance of the policy network during training. The Rank IC assesses the association between the ordinal positions of alpha values and the sequences of stock return rankings. Rank IC is just the IC of ranked data, defined as $RankIC\left(\mathbf{z}^{\prime}_{t},\mathbf{y}_{t}\right)=IC\left(r(\mathbf{z}^{\prime}_{t}),r(\mathbf{y}_{t}\right))$, where $r(\cdot)$ is the ranking operator. This metric is the higher the better. It can be observed that our method can outperform the other RL algorithms in all six indices, especially at the end of the training process. Compared to the best-performing PPO algorithm adopted by AlphaGen, our algorithm improves by 3.83% on the metric, which further validates the effectiveness of discarding elements related to the critic network, introducing a subtractive baseline value, as well as the reasonable reward shaping.
V-C The Studies of the Time-varying Reward Shaping
Figure 6: Performance of QFR with different slope $\eta$, time delay $\alpha$, and maximum IR test value $\delta$, given the same number of iterations.
We study the impact of variations in the slope $\eta$, time delay $\alpha$, and maximum IR test value $\delta$ in reward shaping on the learning process. As shown in Fig. 6, we find that a larger $\eta$ tends to achieve high performance. However, when $\eta$ is too large, the intensity of IR test becomes excessive, causing the policy network to continuously receive discounted rewards after $\alpha$ steps, which harms the efficiency of QFR. Therefore, an appropriate slope of $2.65\times 10^{-6}$ can better trade off between performance and efficiency. The time delay $\alpha$ becomes smaller as the alpha factor quality improves, but an $\alpha$ less than $7\times 10^{4}$ will have a negative impact on the alpha factor quality. We speculate that after the alpha factor quality becomes acceptable, it is appropriate to further optimize the factor quality using IR test mechanism, whereas introducing IR test too early in the training process may not have this effect. An interesting behavior to note is that the training process is highly robust to the maximum IR test value $\delta$. We settled on a value of $0.3$ for this parameter.
V-D Factors Evaluation
@@
We applied the baseline algorithms and our QFR algorithm to the constituent stocks of the CSI300 and CSI500 indices, and evaluated their performance based on two metrics: IC and Rank IC. The IC is delineated in (2). Both of the metrics are the higher the better. Note that all the algorithm is only optimized against the IC metric. The results are shown in Table V. The performance of MLP, XGBoost, and LightGBM is inferior due to the usage of the open-source set of alphas. On the other hand, AlphaGen tends to converge to local optima, leading to overfitting on the training data. Although GP mitigates this issue by preserving a diverse population, it still struggles to generate alphas that exhibit synergy in combination. The QFR algorithm excels in solving MDPs for mining formulaic alpha factors compared to AlphaGen. We attribute the superior performance of QFR to the acceleration of the convergence process by discarding elements related to the critic network, as well as the reasonable reward shaping.
@@ V-E Investment Simulation
TABLE V: Performance of Mined Factors on CSI300 and CSI500
@@
V-F Ablation Study
Figure 9: Ablation experiments in the constituent stocks of CSI300, CSI500, CSI1000, SPX, DJI, and NDX, where the baseline and reward shaping are removed separately. All the curves are averaged over 5 different random seeds, and half of the standard deviation is shown as a shaded region.
To investigate the role of the two improvements of QFR, we designed two variants including: QFR without baseline represents that the policy network is updated directly using REINFORCE, and utilizes the reward function defined in (11); QFR without reward shaping represents that IC is used directly as reward without any IR test. Fig. 9 presents the ablation results of the two improvements. The QFR can provide prominent performance among all six index constituent stocks, indicating that all improvements play important roles. When removing the baseline leads to a degradation in the quality of the factors, it implies that the baseline can help effectively reduce the variance in the training process and achieve better performance. While QFR without reward shaping learns faster during the beginning learning stage, it finally achieves worse performance, which may be due to an overemphasis on the absolute returns of factors while neglecting their volatility resistance, leading to a local optimum trap. In summary, our proposed QFR conditioned on all improvements gives the best performance, indicating that baseline and reward shaping are complementary parts of QFR.
@@ VI Conclusion
In this paper, we have proposed a novel reinforcement learning algorithm, QuantFactor REINFORCE (QFR), for mining formulaic alpha factors. QFR leverages the advantages of discarding the critic network while theoretically addressing its limitations of high variance by introducing a greedy baseline. Additionally, the incorporation of IR as a reward shaping mechanism encourages the generation of stable alpha factors that can better adapt to changing market conditions.